OpenStudy (anonymous):

How many solutions does the system have 6x + 2 = y 3y – 18x = 12 one solution two solutions infinitely many solutions no solution

OpenStudy (shinalcantara):

Problems regarding linear equations with two unknowns can be solved using Substitution and/or Elimination Method. For this let's have elimination. For you to eliminate one variable, 6x + 2 = y -----equation 1 3y - 18x = 12 ----equation 2 ------------------------------ From equation 1: multiply equation 1 by 3 [6x + 2 = y] x3 18x + 6 = 3y Transpose 3y to the left and 6 to the right -3y + 18x = -6 -----equation 1' ------------------------------- Add equations 1' and 2 -3y + 18x = -6 + 3y - 18x = 12 0 + 0 is not equal to 6 ------------------------------- When this occurs to the system of equations, this only means that there's no solution since we can't get the values of x and y and that no values of x and y could be true to both equations

OpenStudy (anonymous):

YOUR THE BEST lol thank you SOOOO much